Solve the systems of equations by graphimg. Then classify the system.
7u+v=48
7u=v+64

Answer by edjones(8007)

7u+v=48
7u=v+64
.
7u+v=48
7u-v=64
--------add
14u=112
.
u=8
v=-8
.
Ed